OpenStudy (pokemon23):

One number is 2 less than another. If 4 times the larger is subtracted from 5 times the smaller the result is 10. Find the numbers One number is 2 less than another. If 4 times the larger is subtracted from 5 times the smaller the result is 10. Find the numbers @Mathematics

OpenStudy (pokemon23):

bath dude

OpenStudy (anonymous):

x = 20 y = 18 from first sentence. x-2 = y that means x has to be the larger number since subtracting from that gets a smaller 5y - 4x = 10 substitute x-2 in for y 5(x-2) -4x =10 5x-10-4x=10 x-10=10 x=20 plug x in the first equation to find y 20-2=y y=18

OpenStudy (pokemon23):

forgot to mention with the same variable